# Rudyak’s conjecture for lower dimensional 1-connected manifolds

### v1

@@ 5. Main result
A degree one map $f:M\to N$ between simply connected $n$-manifolds for $n\leq 8$ satisfies the Rudyak conjecture,
@@
When $n\leq 5$ this is proven in Proposition 3.1. In view of Proposition 3.2 it suffice to show that there is no degree one map $f:M\to N$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\geq 3$.
The case of $n=6$ is taken care of in Corollary 5.3.
We consider the case when $n\geq 7$ and $M$ is spin. Let $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=2$. In view of Proposition 5.5 and Proposition 3.3 we may assume that $f$ induces isomorphisms of homotopy groups of dimension $\leq 3$. Then by Lemma 5.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$.
If $M$ is non-spin and $n\geq 7$, Theorem 4.1 implies that
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq 3.$

### v2

@@ 4. Main result
A degree one map $f:M\to N$ between simply connected spin $n$-manifold for $n\leq 8$ satisfies the Rudyak conjecture,
@@
When $n\leq 5$ the statement is proven in Proposition 3.1. In view of Proposition 3.2 it suffice to show that there is no degree one map $f:M\to N$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\geq 3$. The case of $n=6$ is taken care of in Corollary 4.3.
We consider the case when $n\geq 7$ and $M$ is spin. Let $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=2$. In view of Proposition 4.5 and Proposition 3.3 we may assume that $f$ induces isomorphisms of homotopy groups of dimension $\leq 3$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f\leq 2$. Then by Lemma 4.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. ∎
Remark 4.7.
There are non-spin 7-manifolds with the LS-category 2. We recall that the Wu manifold SU(3)/SO(3) is simply connected and non-spin. A 7-manifold $M$ constructed from the product of a 2-sphere and the Wu manifold
$S^{2}\times(SU(3)/SO(3))$
by a 2-surgery on the factor $S^{2}$ is called a gyration of the Wu manifold [25]. It is simply connected and non-spin. Below we compute its LS-category.
Proposition 4.8.
The LS-category of the gyration of the Wu manifold is 2.
Proof.
Let $Y=SU(3)/SO(3)$. By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}Y=2$. The bordism $X$ defined by the 2-surgery on $Y\times S^{2}$ is homotopy equivalent to the union $(Y\times S^{2})\cup(\{y_{0}\}\times D^{3})$ which is homotopy equivalent to the half-smash product
$Y\rtimes S^{2}=(Y\times S^{2})/(\{y_{0}\}\times D^{3}).$
It is known that for the half-smash product [22]
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(A\rtimes B)=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}A.$
Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}Y=2$. Let $M$ be the gyration of the Wu manifold. Then $X$ is homotopy equivalent to $X^{\prime}=M\cup D^{5}$. The rest of the proof coincides with the end of the proof of Proposition 3.3 (b). Since the inclusion $j:M\to X^{\prime}$ is a 4-equivalence, $\Omega(j):\Omega(M)\to\Omega(X^{\prime})$ is a 3-equivalence. By Proposition 2.5
$\ast^{3}\Omega(j):\ast^{3}\Omega(M)\to\ast^{3}\Omega(X^{\prime})$
is a $(2+2+3=7)$-equivalence. This implies that the map $q:G_{2}(M)\to Z$ to the pull-back in the diagram
[…]
is a $7$-equivalence. Since $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X^{\prime}=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X=2$, there is a section $\sigma:X^{\prime}\to G_{2}(X^{\prime})$ which defines a section $s:M\to Z$ of $p^{\prime}$. Since the map $q$ is a $7$-equivalence, $s$ can be lifted with respect to $q$. Thus, $p_{M}$ admits a section and, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$.
Since $M$ is not homeomorphic to a sphere, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M>1$ and we obtain that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=2$. ∎